OpenStudy (automaticloveletter):

Which of the following ordered pairs represents the solution to the system given below? 2x + y = 20 4x − 2y = 40 (10, 0) (0, 10) (4, 2) (10, −10)

OpenStudy (igreen):

You can use substitution to solve.

OpenStudy (igreen):

\(\sf 4x - 2y = 40\) Subtract 4x to both sides: \(\sf -2y = -4x + 40\) Divide -2 to both sides: \(\sf y = 2x - 20\) Now plug in \(\sf 2x - 20\) into the first equation for \(\sf y\) and solve for \(\sf x\), then plug that into any of the two equations to find \(\sf y\).

OpenStudy (igreen):

Can you do that?

OpenStudy (automaticloveletter):

I'd rather have you walk me through it can you?

OpenStudy (igreen):

Okay, we have: \(\sf 2x + (2x - 20) = 20\) Can you solve that for \(\sf x\)?

OpenStudy (automaticloveletter):

10

OpenStudy (automaticloveletter):

So it's 10, ?

Vocaloid (vocaloid):

continuing from here, we know that x = 10 since you just solved for it let's take one of our original equations 2x + y = 20 let x = 10 and solve for y